# Analysis and Optimized CXL-Attached Memory Allocation for Long-Context LLM Fine-Tuning

### v1

@@ Abstract
The growing prevalence of Large Language Models (LLMs) and their substantial memory requirements have prompted renewed interest in CPU offloading as a method to compensate for limited GPU memory. In particular, when CPU memory is leveraged to temporarily store intermediate states of LLMs, CPU memory becomes a new bottleneck and soon reaches the capacity limitation of commodity CPUs. In this work, we investigate the effectiveness of Compute Express Link (CXL) add-in card (AIC) memory as an extension to CPU memory, enabling larger model sizes and longer context lengths during fine-tuning. Through extensive benchmarking, this study quantifies the performance overhead introduced by transferring data between CXL memory, CPU, and GPUs, focusing on how concurrency and data volume influence bandwidth utilization and latency. This study also compares CPU-based optimizer steps when model parameters, gradients, and optimizer states reside in local memory versus CXL memory, revealing that naive adoption of CXL often degrades performance during the optimizer phase. To overcome these challenges, this study proposes a CXL-aware allocation to strategically partition CPU offloading workloads across both local and CXL memory. This study further demonstrates that employing multiple AICs significantly reduces bandwidth contention, thus improving scalability. Experimental results show that these optimizations enable efficient long-context LLM fine-tuning, underscoring CXL as a promising avenue for unlocking the full potential of CPU offloading in long-context LLM fine-tuning.
Index Terms:
Compute Express Link, Memory Expansion, CPU offloading, Large Language Models, Training
@@ I Introduction
The rapid growth of Large Language Models (LLMs) and the ever-growing number of model parameters have introduced substantial challenges related to memory capacity [1]. As these models frequently exceed available GPU memory, performance bottlenecks arise during both training and deployment. Applications that require longer context lengths [2], such as long chain of thought [3, 4], generative agent [5], in-context learning [6], retrieval-augmented generation [7, 8], and multimodal tasks [9, 10], are experiencing rapid growth. To enhance performance in these scenarios, fine-tuning models on long-context datasets has become increasingly important [11, 12, 13, 14, 15]. However, such fine-tuning introduces substantial memory overhead, primarily due to the need to store intermediate activation values that scale proportionally to the context length [16, 17, 18, 19]. Additionally, limited memory resources restrict the batch size for LLM training, consequently limiting achievable throughput. To address these constraints, especially in resource-limited environments, offloading strategies such as CPU offloading and solid-state drive (SSD) offloading have been proposed [20, 21]. While CPU offloading involves temporarily migrating LLM training or inference states from GPU to CPU (system) memory, SSD offloading further offloads these states onto SSDs, leveraging their larger and more cost-effective storage capacities. However, SSD offloading suffers from inherent performance and endurance limitations associated with NAND flash memory. Consequently, CPU offloading has received significant attention [20, 22, 23, 24]. CPU offloading involves transferring memory components, including model parameters, gradients, optimizer states, and occasionally checkpointed activations, from limited GPU memory to larger-capacity system memory. Although CPU offloading enables the fine-tuning of larger models and supports longer context lengths on GPU-constrained systems, the capacity of the system memory remains a significant challenge, particularly given the continuously growing model sizes and the demand for longer context and larger batch to improve throughput.
@@
System memory capacity is fundamentally limited by CPU specifications (e.g. 192 GB for most consumer grade CPUs [25]), available DIMM slots, and per-module DIMM capacity. To overcome these constraints, Compute Express Link (CXL) technology has emerged as a promising alternative, providing a viable solution to memory bottlenecks encountered during CPU offloading in LLM fine-tuning [26, 27, 28]. Leveraging the PCIe interface and high-performance DRAM, CXL provides a high-bandwidth, low-latency interconnect, enabling expansion of system memory capacity beyond traditional DIMM limitations without incurring the performance penalties of NAND flash-based SSDs. Specifically, CXL Type 3 devices facilitate memory expansion via add-in card (AIC) [29, 30], which are typically recognized by the host operating system as CPU-less Non-Uniform Memory Access (NUMA) nodes. Consequently, applications can access CXL-attached memory similarly to remote DRAM, albeit with distinct performance characteristics.
By utilizing CXL-attached memory, systems can support significantly larger memory footprints, effectively meeting the substantial memory requirements associated with fine-tuning long-context LLMs without resorting to costly high-capacity DIMMs or being constrained by limited DIMM slots [31]. However, our findings demonstrate that naively integrating CXL-attached memory into existing CPU offloading workflows does not inherently ensure optimal performance. Instead, this integration can introduce new performance issues, including data transfer overhead, bandwidth contention, and increased latency due to PCIe-based interconnections. This paper analyzes and strategically mitigates these performance bottlenecks, enabling effective utilization of CXL-attached memory to enhance CPU offloading performance in LLM fine-tuning.
@@
CPU offloading techniques, such as ZeRO-Offload [20], temporarily store model parameters, gradients, and optimizer states from the GPU to the system memory (See Figure 1). While these techniques effectively reduce GPU memory consumption, they introduce frequent data transfers between the GPU and system memory. Additionally, CPU-based operations, like optimizer steps, become susceptible to transfer latency between the CPU and system memory. When integrating CXL-attached memory as an extension or replacement for local DRAM, its distinct performance characteristics, particularly higher latency and differing bandwidth compared to local DIMMs, need to be carefully addressed [32]. Intuitive CXL integrations, including tiered memory system [33, 34] or interleaving approach for CXL-attached memory [35], are typically implemented transparently at the Linux kernel level without requiring application-specific knowledge. Nonetheless, such general-purpose integrations often yield suboptimal performance when running specialized workloads. Consequently, custom CXL memory management policies explicitly tailored to the characteristics of long-context LLM fine-tuning workloads are necessary [36, 37, 38].
Studies have been proposed to explore CXL-attached memory for LLM workloads. For instance, Wang et al. [28] report the end-to-end performance of CPU offloading with CXL-attached memory and Tang et al. [39] leverage CXL-attached memory to store KV cache in LLM inference. Nonetheless, these studies primarily focus on characterizing general CXL-attached memory performance without analyzing details and proposing optimizations tailored specifically to CPU offloading workloads, resulting in a significant gap in understanding complex performance interactions such as frequent GPU-CPU data transfers, latency-sensitive CPU-based optimizer steps, and multi-GPU bandwidth contention on shared CXL AICs during long-context fine-tuning. To address these challenges, this study first performs benchmarking to quantify the performance overhead associated with utilizing CXL-attached memory in long-context CPU offloading scenarios. The analysis identifies two primary performance bottlenecks. The first bottleneck arises from significant performance degradation during CPU-based optimizer operations when model parameters, gradients, or optimizer states reside in higher latency CXL-attached memory rather than in local DRAM. The second bottleneck emerges from bandwidth contention when multiple GPUs simultaneously access the same CXL AIC, which limits overall system scalability.
To mitigate these bottlenecks, this study proposes a CXL-aware long-context LLM fine-tuning comprising two key components: (1) CXL-aware Memory Allocation, which strategically allocates offloaded data (model parameters, gradients, optimizer states) and activation values by considering their access patterns and latency sensitivity, thereby prioritizing memory placement according to specific workload requirements; and (2) Multi-AIC Striping, which allow CPU to leverage aggregated bandwidth of multiple CXL AICs and also alleviates bandwidth contention in multi-GPU system by distributing memory allocations across multiple CXL AICs. The multi-AIC striping method increases the available aggregate bandwidth and significantly reduces bottlenecks during intensive data transfer phases. Through these optimizations, this study demonstrates that CXL-attached memory can effectively enhance CPU offloading capabilities for long-context LLM fine-tuning, enabling training of larger models and handling longer contexts on resource-constrained hardware, while achieving performance comparable to configurations utilizing exclusively local DRAM. The main contributions are as follows.
@@
    To the best of our knowledge, this paper presents the first detailed empirical study on the performance of CXL AIC memory specifically for long-context LLM fine-tuning with CPU offloading.
@@
    This study identifies and analyzes key performance bottlenecks introduced by naive CXL adoption.
@@
    The CXL-aware memory allocation is proposed to efficiently manage local DRAM and CXL-attached memory based on workload characteristics.
@@
    The multi-AIC striping mitigates bandwidth issues in multi-GPU setups and improves scalability.
  5. 5.
    Experimental results demonstrate that CXL-attached memory can sustain high performance, establishing CXL as a promising solution for next-generation AI workloads.
@@ II Background
Figure 1: Example of long-context CPU offloading with activation checkpointing with a transformer model composed of 4 transformer blocks. Arrows indicate data transfers over PCIe: $P_{i}$ represent model parameters (e.g., attention projection parameters, feed-forward network parameters) for a specific block. $A_{i}$ represent checkpointed input activations for block. $G_{i}$ represent gradients corresponding to the parameters of block. The numbered steps illustrate the data movement and computation flow.
@@ II-A CPU Offloading for Long-Context Fine-Tuning
ZeRO-Offload [20] is a widely used technique to train LLMs on systems with limited GPU memory. It conserves GPU resources by transferring model parameters, gradients, and optimizer states from GPU to system memory, and only retrieves them back to GPU memory when required for computation. To further reduce memory usage, ZeRO-Offload can be combined with techniques such as Flash-Attention [40, 41], Liger-Kernel [42], and gradient checkpointing (activation checkpointing) [17]. Flash-Attention efficiently computes attention without fully materializing the attention matrix, ensuring that peak memory scales linearly rather than quadratically with context length. Liger-Kernel optimizes large intermediate tensor usage during cross-entropy calculation by employing a FusedLinearCrossEntropy mechanism. Notably, the intermediate tensor usage also scales with context length and vocabulary size. Activation checkpointing reduces peak memory by storing only a subset of activations during the forward pass and recomputing them during the backward pass. Although these techniques significantly reduce GPU memory usage, they are still insufficient in long-context fine-tuning scenarios. This is because the size of checkpointed activations grows substantially with context length, making it necessary to offload them to system memory and retrieve them when needed [19]. Since model parameters, gradients, optimizer states, and checkpointed activations are all stored in system memory, the demand on system memory increases rapidly and can become a major bottleneck.
@@
Figure 1 illustrates the aforementioned integrated approach, referred to as long-context CPU offloading with offloaded activation checkpointing, using a transformer model with four transformer blocks as an example. Each transformer block contains parameters for attention projections and feedforward layers. The workflow operates as follows: (1) First, necessary parameters are loaded from CPU to GPU memory on a tensor-by-tensor basis. (2) Next, the GPU performs forward computations using these parameters. (3) Checkpointed activations for each transformer block are offloaded to CPU memory. (4) Once the forward pass concludes, the backward pass requires parameters and previously checkpointed activations. (5) These data are then reloaded onto the GPU, which recomputes necessary activations to perform backpropagation. (6) Gradients computed on the GPU are subsequently offloaded to CPU memory, (7) enabling optimizer updates (e.g., using Adam) to execute entirely on the CPU after completing the backward pass. During optimizer steps, full precision parameters, optimizer states, and gradients reside primarily in CPU memory. Such an approach minimizes the volume of data transferred between the GPU and the CPU during each training iteration.
@@
To clarify the memory footprint of the workflow shown in Figure 1, the GPU memory usage is first examined. During CPU offloading, the GPU is dedicated solely to computation and retains minimal data: model parameters are streamed block by block, and the corresponding activations and gradients are kept only until each block’s computation is complete. These are then immediately offloaded to system memory or discarded, shifting the primary memory burden to the system memory. On the other hand, the system memory usage is detailed in Table I. The upper half lists components frequently transferred between CPU and GPU during forward and backward passes, while the lower half lists components stored on the CPU for optimizer updates. The memory usage for model parameters and gradients depends on their precision: bf16 requires 2 bytes per parameter, while fp32 requires 4 bytes per parameter. For the Adam optimizer, the optimizer states (momentum and variance) require $8\times P$ bytes in fp32, doubling the memory of gradients due to maintaining two states per parameter. For checkpointed activations, each GPU stores unique activations, scaled by $N_{g}$. Checkpoints are saved for each transformer block’s input, totaling $L$ blocks, with each activation sized at $B\times C\times H$ elements, stored in bf16 (2 bytes per element).
@@
TABLE I: Breakdown of system memory components during CPU offloading. $P$: total parameters; $N_{g}$: number of GPUs; $B$: batch size per GPU; $C$: context length; $L$: number of transformer blocks; $H$: hidden size.
@@
II-B CPU Memory Bottleneck under Long-Context Offloading
@@
In long-context CPU offloading scenarios, memory demand shifts predominantly from GPU memory to system memory. As a result, system memory capacity becomes a key factor in determining the feasible model size and the maximum context length. A larger system memory allows training with larger models, supports longer context windows, and enables sufficiently large batch sizes to achieve optimal performance. To illustrate this behavior, a motivational experiment was conducted using the 12B model to measure memory requirements and throughput across different context lengths and batch sizes in a 2-GPU setting (hardware specifications are listed in Table II). In the first part of the experiment, the batch size is fixed at 5, and the context length is varied from 512 to 32K tokens. The choice of 32K is based on previous long-context fine-tuning studies [11, 12, 13, 14, 15], which commonly use datasets with context lengths around 32K. For example, LongAlpaca [11] ranges from 3K to 9K, FILM [12] spans 4K to 32K, LongWriter [14] ranges from 2K to 32K, and LongAlign [15] ranges from 8K to 64K, with 90% of samples below 32K. In the second part, the context length is fixed at 4K, while the batch size is varied from 1 to 48 to observe changes in throughput and memory usage. The results are presented in Figures 4 and 4.

### v2

@@ Abstract
The substantial memory requirements of Large Language Models (LLMs), particularly for long-context fine-tuning, have renewed interest in CPU offloading to augment limited GPU memory. However, as context lengths grow, relying on CPU memory for intermediate states introduces a significant bottleneck that can exhaust the capacity of mainstream client platforms. To address this limitation, this work investigates the effectiveness of Compute Express Link (CXL) add-in card (AIC) memory as an extension to CPU memory, enabling larger model sizes and longer context lengths during fine-tuning. Extensive benchmarking reveals two critical challenges. First, current deep learning frameworks such as PyTorch lack fine-grained, per-tensor control over NUMA memory allocation, exposing only coarse, process-level policies. Second, due to this lack of control, when the memory footprint of fine-tuning is offloaded across local DRAM and CXL-attached memory, naively placing optimizer data in higher-latency CXL leads to substantial slowdowns in the optimizer step (e.g., $\sim$4× once data exceeds $\sim$20M elements). To overcome these challenges, this work introduces a PyTorch extension that enables tensor-level system memory control and a CXL-aware memory allocator that pins latency‑critical tensors in local DRAM while maximizing bandwidth by striping latency‑tolerant tensors across one or more CXL devices. Evaluated on a real hardware setup with 7B and 12B models, 4K–32K contexts, and a single GPU, our approach recovers throughput to 97–99% of DRAM‑only with a single AIC and $\approx$100% with two AICs, delivering up to 21% improvement over naive interleaving while preserving DRAM‑like DMA bandwidth for GPU transfers. These results show that carefully managed CXL‑attached memory is a practical path to scaling long‑context fine‑tuning beyond DRAM limits.
@@ I Introduction
The rapid growth of Large Language Models (LLMs) and their ever-increasing parameter counts have introduced significant challenges in memory capacity [1]. As these models frequently exceed available GPU memory, performance bottlenecks arise during both training and deployment. Memory requirements are further exacerbated by the push toward longer context lengths [2], which arise from applications such as long chain-of-thought reasoning [3, 4], generative agents [5], in-context learning [6], retrieval-augmented generation [7], and multimodal tasks [8], all of which are experiencing rapid growth. To support these scenarios, fine-tuning LLMs on long-context datasets has become increasingly important [9, 10, 11, 12, 13]. However, long-context fine-tuning imposes substantial memory overhead, primarily from storing intermediate activation values that scale with context length [14, 15]. Furthermore, limited memory resources restrict batch size during training, thereby constraining throughput.
To address these constraints, particularly in resource-limited environments, offloading strategies, such as CPU offloading and solid-state drive (SSD) offloading, have been proposed [16, 17]. CPU offloading migrates model states, such as parameters, gradients, optimizer data, and occasionally checkpointed activations, from GPU memory to system memory (see Figure 1), while SSD offloading further offloads these states onto SSDs, leveraging the larger and more cost-effective capacity of SSDs. However, SSD offloading suffers from inherent performance and endurance limitations of NAND flash memory, making CPU offloading the more practical and widely adopted approach [16, 18, 19, 20]. Although CPU offloading enables the fine-tuning of larger models and supports longer context lengths on GPU-memory-constrained systems, system memory capacity itself become the bottleneck instead, especially as model sizes continue to grow and the demand for longer contexts and larger batch sizes increases.
@@
System memory capacity on mainstream client platforms (often  192–256 GB today [21]) is constrained by CPU/chipset limits, DIMM slots, and module density. To overcome these constraints, Compute Express Link (CXL) technology has emerged as a promising alternative, providing a viable solution to memory bottlenecks encountered during CPU offloaded long-context LLM fine-tuning [22, 23, 24]. Leveraging PCIe and DRAM, CXL-attached memory expands host capacity beyond DIMM density/slot limits without the performance penalties of NAND-flash SSDs or the cost of high-capacity DIMMs [25]. In particular, CXL Type-3 add-in cards (AICs) provide memory expansion [26, 25] that the host OS typically exposes as CPU-less NUMA nodes, allowing applications to access them similarly to remote DRAM, though with distinct performance characteristics. This capacity enables long-context LLM fine-tuning without being limited by the system memory size; however, since CPU offloading workloads are sensitive to system memory access latency, naively integrating CXL-attached memory into existing CPU offloading workflows does not inherently ensure optimal performance.
Consequently, custom CXL memory management policies tailored to offloading workloads are required [27, 28]. Recent work has begun exploring CXL-attached memory for LLM workloads. For example, Wang et al. [24] evaluate end-to-end performance of CPU offloading with CXL, while Tang et al. [29] employ CXL memory to store the KV cache during inference. However, these studies mainly characterize general performance without analyzing workload-specific behavior or proposing optimizations tailored to CPU offloaded long-context LLM fine-tuning tasks. This leaves an open gap in understanding interactions such as frequent GPU–CPU transfers and latency-sensitive optimizer phases during long-context fine-tuning.
@@
In practice, CPU offloading techniques such as ZeRO-Offload [16] reduce GPU memory consumption by transferring model parameters, gradients, and optimizer states to system memory. While this effectively alleviates GPU pressure, it introduces frequent data transfers between GPU and CPU, and exposes performance sensitivity to memory access latency. When CXL-attached memory is used as an extension or replacement for local DRAM in offloading scenarios, its distinct performance characteristics, including higher latency and lower bandwidth compared to local DIMMs, need to be carefully considered [30]. Generic operating system (OS)-level mechanisms, such as tiered memory systems [31, 32] or interleaving policies [33], provide transparent support but often yield suboptimal performance for specialized workloads. To clarify these implications, this work benchmarks CXL-attached memory under long-context CPU offloading workloads and identifies two primary performance challenges and one key performance characteristic.
First, current deep learning frameworks such as PyTorch [34] enforce a single uniform allocation policy across DRAM and CXL-attached memory. Because memory is managed only at the process level, users are limited to coarse-grained tools like numactl [35], making fine-grained, per-tensor placement impossible and often leading to latency-sensitive tensors being allocated in CXL-attached memory. Second, CPU-based optimizer phases are particularly latency sensitive: read–modify–write loops over parameters, gradients, and optimizer states degrade significantly if data reside in CXL-attached memory rather than DRAM, as the memory accesses of the CPU-based optimizer step dominate the critical path and directly reduce throughput. On the other hand, this work identifies a key performance characteristic: in the GPU pipeline, asynchronous DMA overlaps data movement with kernel execution [^1] A kernel is a low-level function that performs a specific tensor operation (e.g., convolution, matrix multiply). Kernel execution is running that function on hardware (CPU/GPU) to carry out the computation in parallel., and because both DRAM and CXL-attached memory traverse the same PCIe path, host-to-device transfer throughput is broadly comparable from either source.
To address these challenges and build on insights from prior analysis, this work introduces two primary optimizations. First, a fine-grained memory controller is designed and implemented as a PyTorch extension to enable per-tensor control over NUMA memory placement. Second, this paper introduces a CXL-aware memory allocator, implemented as a greedy policy that leverages the fine-grained memory controller to partition tensors by latency sensitivity. Latency-critical data are placed in local DRAM, while latency-tolerant data are directed to CXL-attached memory, with interleaving applied when beneficial to maximize bandwidth. In other words, through the introduced components, latency-bound optimizer states/updates in CPU-offloaded fine-tuning are pinned in DRAM, while bandwidth-bound GPU transfers are striped over CXL AICs via tensor-level placement. Together, these optimizations demonstrate that CXL-attached memory can effectively expand capacity for long-context LLM fine-tuning while achieving performance nearly identical to DRAM-only configurations. Across 7B/12B models and 4K–32K contexts with a single GPU, our CXL‑aware memory allocator restores single‑AIC throughput to 97–99% of DRAM‑only and matches DRAM‑only with dual AICs, outperforming naive interleaving by up to 21%. The main contributions of this study are as follows.
@@
    To the best of our knowledge, this paper presents the first empirical characterization of CXL-attached memory for long-context LLM fine-tuning, identifying and analyzing the key performance bottlenecks (e.g., optimizer latency sensitivity) introduced by naive CXL adoption.
@@
    A PyTorch extension is implemented to enable precise, per-tensor memory placement on specified NUMA nodes, a crucial capability for heterogeneous memory systems.
@@
    A CXL-aware memory allocator is introduced as a greedy policy that maps tensors by latency sensitivity and leverages interleaving when beneficial to maximize bandwidth and minimize latency-induced slowdowns.
@@
    Real-world experimental results on CXL devices demonstrate that CXL-attached memory can expand capacity while delivering performance comparable to DRAM-only baselines for long-context LLM fine-tuning workloads.
@@ II-A CPU Offloading for Long-Context Fine-Tuning
ZeRO-Offload [16] is a widely used technique to train LLMs on systems with limited GPU memory. It conserves GPU resources by transferring model parameters, gradients, and optimizer states from GPU memory to system memory, and only retrieves them back to GPU memory when required for computation. To further reduce memory usage, ZeRO-Offload can be combined with techniques such as Flash-Attention [36], Liger-Kernel [37], and gradient checkpointing (activation checkpointing) [15]. Flash-Attention efficiently computes attention without fully materializing the attention matrix, ensuring that peak memory scales linearly rather than quadratically with context length. Liger-Kernel optimizes large intermediate tensor usage during cross-entropy calculation by employing a FusedLinearCrossEntropy mechanism. Notably, the intermediate tensor usage also scales with context length and vocabulary size. Activation checkpointing reduces peak memory by storing only a subset of activations during the forward pass and recomputing them during the backward pass. As context length grows, the volume of checkpointed activations increases, necessitating offload to system memory and on-demand retrieval [15].
Figure 1: Example of long-context CPU offloading with activation checkpointing with a transformer model composed of 4 transformer blocks. Arrows indicate data transfers over PCIe: $P_{i}$ represent model parameters (e.g., attention projection parameters, feed-forward network parameters) for a specific block. $A_{i}$ represents checkpointed input activations for the block. $G_{i}$ represents gradients corresponding to the parameters of the block. The numbered steps illustrate the data movement and computation flow.
@@
Figure 1 illustrates the aforementioned integrated approach, referred to as long-context CPU offloading with offloaded activation checkpointing, using a transformer model with four transformer blocks as an example. Each transformer block contains parameters for attention projections and feedforward layers. The workflow operates as follows: (1) First, necessary parameters are loaded from CPU to GPU memory on a tensor-by-tensor basis. (2) Next, the GPU performs forward computations using these parameters. (3) Checkpointed activations for each transformer block are offloaded to CPU memory. (4) Once the forward pass concludes, the backward pass requires parameters and previously checkpointed activations. (5) These data are then reloaded onto the GPU, which recomputes necessary activations to perform backpropagation. (6) Gradients computed on the GPU are subsequently offloaded to CPU memory, (7) enabling optimizer updates (e.g., using Adam) to execute entirely on the CPU after completing the backward pass. During optimizer steps, full precision parameters, optimizer states, and gradients reside primarily in CPU memory. Such an approach minimizes the volume of data transferred between the GPU and the CPU during each training iteration. Notably, the aforementioned workflow, which combines ZeRo-offload, Flash-Attention, Kiger-Kernel, and offloaded activation checkpointing, is considered the baseline of this study.
@@
To clarify the memory footprint of the workflow shown in Figure 1, the GPU memory usage is first examined. During CPU offloading, the GPU is dedicated solely to computation and retains minimal data: model parameters are streamed block by block, and the corresponding activations and gradients are kept only until each block’s computation is complete. These are then immediately offloaded to system memory or discarded, shifting the primary memory burden to the system memory. On the other hand, the system memory usage is detailed in Table I. The upper half lists components frequently transferred between CPU and GPU during forward and backward passes, while the lower half lists components stored on the CPU for optimizer updates. The memory usage for model parameters and gradients depends on their precision: bf16 requires 2 bytes per parameter, while fp32 requires 4 bytes per parameter. Notably, Zero-Offload uses bf16 on GPUs to manage the huge memory footprint of activations and maximize throughput, while strategically using fp32 on CPU for the sensitive optimizer calculations to ensure the model learns correctly and stably. For checkpointed activations, each GPU requires a unique set of activations; thus, the total system memory usage for these activations is scaled by $N_{g}$. Checkpoints are saved for each transformer block’s input, totaling $L$ blocks, with each activation sized at $B\times C\times H$ elements, stored in bf16 (2 bytes per element). For the Adam optimizer, the optimizer states (momentum and variance) require $8\times P$ bytes in fp32, doubling the memory of gradients due to maintaining two states per parameter. Despite the aforementioned workflow having substantially reduced GPU memory usage, the workflow remains insufficient for long-context fine-tuning. This is because the memory required for activations grows enormously with sequence length. As a result, with parameters, gradients, optimizer states, and these massive activations all resident in system memory, memory pressure escalates rapidly, and system memory itself becomes the dominant bottleneck.
@@
TABLE I: Breakdown of system memory components during CPU offloading. The upper half depicts GPU-CPU transfer size, and the Lower half depicts system memory usage. $P$: total parameters; $N_{g}$: number of GPUs; $B$: batch size per GPU; $C$: context length; $L$: number of transformer blocks; $H$: hidden size.
@@
II-B Memory Bottleneck under Long-Context Offloading
In the demanding scenario of training long-context LLMs with CPU offloading techniques, memory demand shifts predominantly from GPU memory to system memory. As a result, system memory capacity becomes a key factor, directly determining the feasible model size, maximum context length, and the batch sizes required to achieve optimal performance. To illustrate this behavior, a motivational experiment was conducted using the 12B model to measure memory requirements and throughput across different context lengths and batch sizes in a 2-GPU setting (hardware specifications are listed in Table II). In the first part of the experiment, the batch size is fixed at 5, and the context length is varied from 512 to 32K tokens. The choice of 32K is based on previous long-context fine-tuning studies [9, 10, 11, 12, 13], which commonly use datasets with context lengths around 32K. For example, LongAlpaca [9] ranges from 3K to 9K, FILM [10] spans 4K to 32K, LongWriter [12] ranges from 2K to 32K, and LongAlign [13] ranges from 8K to 64K, with 90% of samples below 32K. In the second part, the context length is fixed at 4K, while the batch size is varied from 1 to 48 to observe changes in throughput and memory usage. The results are presented in Figures 4 and 4.